# Coefficient Shape Transfer Learning for Functional Linear Regression

### v2

@@ 3.1 Identification through Regularization
[…]
@@
A grid search over $(\lambda,\tilde{\lambda})$ yields a sequence of candidate informative sets for $\widehat{\mathcal{A}}$. Each candidate is evaluated via a bootstrap procedure: the target coefficient $\beta^{(0)}(t)$ is estimated using transfer learning, residuals $\hat{\epsilon}_{n}=y^{(0)}_{n}-\hat{y}^{(0)}_{n}$ are computed, and bootstrap responses $\tilde{y}^{(0)}_{n}=\hat{y}^{(0)}_{n}+\hat{\epsilon}^{*}_{n}$ are generated, where $\hat{\epsilon}^{*}_{n}$ is obtained by resampling $\hat{\epsilon}_{n}$ with replacement. The model is then refitted to obtain a bootstrap estimate, and the mean squared prediction error (MSE) is calculated. This process is repeated (e.g., 1000 times) to compute the average MSE for each candidate set. The final informative set $\widehat{\mathcal{A}}$ is selected as the one minimizing the average MSE. After identifying the informative set, a post-selection procedure can be applied to remove source domains with significantly slower spectral decay rates and small coefficient magnitudes.
@@ 4.1 General Setting
The coefficient function of the target model is determined by $B_{1}$. To investigate the impact of non-informative source domains, we designate the $6$-th to $9$-th source domains as non-informative. The oracle informative set consists of the $1$-st to $5$-th source domains.
@@ 5.2 Data Pre-processing
We first remove the outliers based on the $\ell^{2}$-norm of the raw activity curves for the ten groups. The outliers are defined as those beyond the interval $[Q_{1}-1.5\times\mbox{IQR},Q_{3}+1.5\times\mbox{IQR}]$. where $Q_{1},Q_{3}$ are the first, third quantile of the function $\ell^{2}$-norm, and $\mbox{IQR}=Q_{3}-Q_{1}$. We transform the raw activity functions into occupation time functions. Occupation time refers to the duration a stochastic process spends within a particular set, such as exceeding a threshold level “$a$”. Let $\{X(t)\colon t\in[0,T]\}$ be a measurable stochastic process representing the complete observed trajectory of a raw activity curve for a given subject. The occupation time that $X(t)$ spends above the threshold $a$ is defined as: $L(a)=\text{Leb}\left(\{t\in[0,T]\colon X(t)>a\}\right)$, where Leb denotes the Lebesgue measure on the real line. The index “$a$” varies over the range of activity levels being studied. As a function of “$a$”, $L(a)$ is bounded between 0 and $T$, is monotonically decreasing, and is right-continuous by definition. The original occupation time functions for the ten groups are displayed in Figure 4. Then we take the square root transformation for the occupation functions and log-transformation for chronological age to stabilize variance. The log-transformed age and the square-rooted occupation time functions are then used as response and covariate variables respectively.
@@ 5.3 Estimation Performance and Comparison
[…]
@@ 6 Conclusions
We propose a novel transfer learning framework to improve the estimation of functional linear models by transferring knowledge of the coefficient shape from source domains to the target domain. Compared to traditional assumptions of coefficient homogeneity, the notion of coefficient shape homogeneity is more inclusive and robust, particularly to covariate scaling. This flexibility enables the proposed method to effectively incorporate relevant structural information from multiple sources, enhancing the accuracy of model estimation in the target domain.
The proposed approach proceeds in two main stages. First, we estimate a shared coefficient shape by aggregating information from both the target domain and source domains. In the second stage, we refine this estimated template by multiplying it with a scale parameter to obtain the final coefficient estimate. We rigorously study the convergence properties of the resulting transfer learning estimator under both non-random and random designs. Theoretical analysis reveals that the convergence rate is influenced by the sample sizes of different domains, the spectral decay rate of covariate covariance operators, the magnitudes of coefficients, and the shape divergence between source domains and the target domain coefficients.
To address the challenge of unknown informativeness, we also introduce a source domain selection procedure that identifies informative source domains without relying on prior knowledge. The practical effectiveness of our method is demonstrated through extensive simulation studies and real data analysis using NHANES physical activity data.
The proposed framework, grounded in the concept of coefficient shape homogeneity, provides a flexible foundation for broader applications and can be extended to other functional linear models, including functional autoregressive models and functional factor models. These extensions are nontrivial and will be pursued in future work.

### v3

@@ 3.1 Identification through Regularization
where $J_{\lambda}(\cdot)$ is a concave penalty function such as the smoothly clipped absolute deviation penalty $J_{\lambda}(x)=\int_{0}^{|x|}\lambda\min\{1,\frac{(\gamma\lambda-t)_{+}}{(\gamma-1)\lambda}\}dt$ (SCAD, see Fan and Li [12]) and minimax concave penalty $J_{\lambda}(x)=\min\{\lambda|x|-{x^{2}}/{2\gamma},\gamma\lambda^{2}/2\}$ (MCP, see Zhang [35]). The concavity property is important for guaranteeing the consistency of informative set identification (see also Jiao and Chan [19], Ma and Huang [24] and Shen and Huang [29]). The penalty term incorporates all the coefficient shape misalignments between the source domains and the target domain and shrinks insignificant coefficient shape misalignments, aiding in the identification of the informative set.
@@
[…]
@@ 4.1 General Setting
The coefficient function of the target model is determined by $B_{1}$. To investigate the impact of non-informative source domains, we designate the 66-th to 99-th source domains as non-informative. The oracle informative set consists of the 11-st to 55-th source domains.
@@ 5.2 Data Pre-processing
We first remove the outliers based on the $\ell^{2}$-norm of the raw activity curves for the ten groups. The outliers are defined as those beyond the interval $Q_{1}-1.5\times\mbox{IQR},Q_{3}+1.5\times\mbox{IQR}]$. where $Q_{1},Q_{3}$ are the first, third quantile of the function $\ell^{2}$-norm, and $\mbox{IQR}=Q_{3}-Q_{1}$. We transform the raw activity functions into occupation time functions. Occupation time refers to the duration a stochastic process spends within a particular set, such as exceeding a threshold level “aa”. Let $\{X(t)\colon t\in[0,T]\}$ be a measurable stochastic process representing the complete observed trajectory of a raw activity curve for a given subject. The occupation time that $X(t)$ spends above the threshold $a$ is defined as: $L(a)=\text{Leb}\left(\{t\in[0,T]\colon X(t)>a\}\right)$, where Leb denotes the Lebesgue measure on the real line. The index “aa” varies over the range of activity levels being studied. As a function of “aa”, $L(a)$ is bounded between 0 and $T$, is monotonically decreasing, and is right-continuous by definition. The original occupation time functions for the ten groups are displayed in Figure 4. Then we take the square root transformation for the occupation functions and log-transformation for chronological age to stabilize variance. The log-transformed age and the square-rooted occupation time functions are then used as response and covariate variables respectively.
@@ 5.3 Estimation Performance and Comparison
1:Fit the functional linear model and obtain residuals: $\widehat{Y}_{i}=\int X_{i}(t)\hat{\beta}^{(0)}(t)dt$, $\hat{\epsilon}_{i}=Y_{i}-\hat{Y}_{i}$.
@@ 6 Conclusions
We have developed a novel transfer learning framework for functional linear models that leverages coefficient shape similarity to transfer structural information across domains. By shifting the focus from strict coefficient homogeneity to coefficient shape homogeneity, our framework substantially broadens the scope of transferable knowledge, offering a more robust and interpretable foundation for cross-domain inference. This flexibility enables the integration of heterogeneous yet structurally related sources, thereby improving estimation accuracy in the target domain.
The proposed approach proceeds in two main stages. First, we estimate a shared coefficient shape by aggregating information from both the target domain and source domains. In the second stage, we refine this estimated template by multiplying it with a scale parameter to obtain the final estimator of the target coefficient. We rigorously study the convergence properties of the resulting final estimator under both non-random and random designs. Theoretical analysis reveals that the convergence rate is influenced by the sample sizes of different domains, the spectral decay rate of covariate covariance operators, the magnitudes of coefficients, and the shape divergence between the source and target domain coefficients. To address the challenge of unknown informativeness, we develop a source domain selection procedure that identifies informative source domains without relying on prior knowledge. The practical effectiveness of our method is demonstrated through extensive simulation studies and real data analysis using NHANES physical activity data. The proposed framework, grounded in the concept of coefficient shape homogeneity, provides a flexible foundation for broader applications and can be extended to other functional linear models, including functional autoregressive models and functional factor models. These extensions are nontrivial and will be pursued in future work.